I can get every part except the last part my answer does not come out right I need to know how to get the final answer.
Find an equation of the line containing the given pair of points (-5,-8), (-1,-2)

Answer by CharlesG2(834)

I can get every part except the last part my answer does not come out right I need to know how to get the final answer.
Find an equation of the line containing the given pair of points (-5,-8), (-1,-2)
m = slope = rise/run = (y2 - y1)/(x2 - x1)
m = (-2 - -8)/(-1 - -5)
m = (-2 + 8)/(-1 + 5)
m = 6/4 = 3/2
the equation of a line is in form y=mx+b, where m is the slope,
and b is the y-intercept (vertical-intercept or point (0,b))
y = 3/2 * x + b (we need to solve for b)
-2 = 3/2 * -1 + b (plugged in (-1,-2)
-2 = -3/2 + b
-1/2 = b
-8 = 3/2 * -5 + b (plugged in (-5,-8))
-8 = -15/2 + b
-1/2 = b again
y = (3/2)x - (1/2) (equation of the line)